OpenStudy (anonymous):

wut is (2 + 2 - 2 / 2 x 2) + 99^10 / (2 + 2 - 2 / 2 x 2) / 4

OpenStudy (anonymous):

nvm its 1.1304776e+19

OpenStudy (amoodarya):

\[\\(2+2-\frac{ 2 }{ 2x^2 })+\frac{ 99^{10} }{\frac{ (2+2-\frac{ 2 }{ 2x^2 }) }{ 4 } }\] do you mean this ?

OpenStudy (anonymous):

yeah, idk how 2 do that

OpenStudy (amoodarya):

\[(4-\frac{ 1 }{ x^2 })+\frac{ 99^{10} }{\frac{ (4-\frac{ 1 }{ x^2 }) }{ 4 } }\\\] \[(4-\frac{ 1 }{ x^2 })+4\frac{ 99^{10} }{\frac{ (4-\frac{ 1 }{ x^2 }) }{ 1 } }\\\]